# Adaptive Mix for Semi-Supervised Medical Image Segmentation

### v1

@@ Abstract
Mix-up is a key technique for consistency regularization-based semi-supervised learning methods, generating strong-perturbed samples for strong-weak pseudo-supervision. Existing mix-up operations are performed either randomly or with predefined rules, such as replacing low-confidence patches with high-confidence ones. The former lacks control over the perturbation degree, leading to overfitting on randomly perturbed samples, while the latter tends to generate images with trivial perturbations, both of which limit the effectiveness of consistency learning. This paper aims to answer the following question: How can image mix-up perturbation be adaptively performed during training? To this end, we propose an Adaptive Mix algorithm (AdaMix) for image mix-up in a self-paced learning manner. Given that, in general, a model’s performance gradually improves during training, AdaMix is equipped with a self-paced curriculum that, in the initial training stage, provides relatively simple perturbed samples and then gradually increases the difficulty of perturbed images by adaptively controlling the perturbation degree based on the model’s learning state estimated by a self-paced regularize. We develop three frameworks with our AdaMix, i.e., AdaMix-ST, AdaMix-MT, and AdaMix-CT, for semi-supervised medical image segmentation. Extensive experiments on three public datasets, including both 2D and 3D modalities, show that the proposed frameworks are capable of achieving superior performance. For example, compared with the state-of-the-art, AdaMix-CT achieves relative improvements of 2.62% in Dice similarity coefficient and 48.25% in average surface distance on the ACDC dataset with 10% labeled data. The results demonstrate that mix-up operations with dynamically adjusted perturbation strength based on the segmentation model’s state can significantly enhance the effectiveness of consistency regularization. The code will be released at https://github.com/Senyh/AdaMix.
@@ 1 Introduction
Medical image segmentation for delineating tissues, organs, and lesions from different imaging modalities is critical for computer-aided clinical applications, such as prognosis and preoperative planning. Deep learning has greatly advanced the progress of medical image segmentation, relying on its powerful representation capability and the availability of large amounts of labeled data (Ronneberger et al., 2015; Milletari et al., 2016). However, obtaining pixel-level annotations is laborious and expensive in medical images, as it requires expert knowledge. Semi-supervised learning (SSL) has the potential to alleviate the label scarcity problem by leveraging a large amount of unlabeled data (Van Engelen and Hoos, 2020).
@@
Figure 1: Comparison among CutMix, UMix (Shen et al., 2023), and our AdaMix. (a) Original image and auxiliary image; (b) CutMix randomly replaces patches of the original image with patches from the auxiliary image, with uncontrollable perturbation degree, leading to an oscillating and unstable consistency regularization process; (c) UMix replaces low-confidence patches in the original image with high-confidence ones from the auxiliary image, often generating perturbed examples with trivial perturbations, thereby limiting the effect of consistency regularization; (d) I-UMix mixes low-confidence patches in the original image with high-confidence ones from the auxiliary image; (e) Our AdaMix synthesizes images with more high-confidence regions in the initial training stage and then gradually increases the perturbation degree to enhance the complexity of the perturbed images as training progresses; (f) The unsupervised learning loss curves; (g) The validation loss curves (and the corresponding Dice Similarity Coefficients on the ACDC test set).
@@
Consistency regularization, one of the most widely-used SSL paradigms, encourages model predictions to be invariant to input perturbations, aiming to sufficiently explore supervision information from unlabeled data (Rasmus et al., 2015; Laine and Aila, 2016; Tarvainen and Valpola, 2017). Building upon the idea of consistency regularization, strong-weak pseudo supervision (Sohn et al., 2020) aims to enforce consistency between the predicted pseudo labels of strongly and weakly perturbed unlabeled images, thereby extracting additional supervisory signals from unlabeled data (Sohn et al., 2020; Chen et al., 2023, 2021; Wu et al., 2022; Shen et al., 2023; Wang et al., 2023b). The key challenge of strong-weak pseudo supervision is determining appropriate strong-perturbation strategies for generating strong-perturbed images. Improper perturbations cannot match the training model’s ability to obtain supervision signals from unlabeled data, misleading the training process (Shen et al., 2023). Recently, CutMix [^1] CutMix is formulated as $\bar{x}=x_{o}\times m+x_{a}\times(1-m)$ and $\bar{y}=y_{o}\times m+y_{a}\times(1-m)$, where $x_{o}$ / $x_{a}$ denotes an original/auxiliary image, $y_{o}$ / $y_{a}$ refers to the corresponding (pseudo) label, and $m$ is a randomly generated mask. $\bar{x}$ / $\bar{y}$ represents the mixed image / (pseudo) label.(Yun et al., 2019), incorporated with image color space transformation, has been widely used to form the strong perturbations (French et al., 2019; Chen et al., 2021; Yang et al., 2022, 2023). As illustrated in Fig. 1(b), CutMix randomly mixes image patches between an original image and an auxiliary image to synthesize a strong-perturbed example. However, such a random mix operation with uncontrollable perturbation degree leads to an oscillating and unstable consistency regularization process [Fig. 1(f)], where the model is prone to over-fit to the randomly perturbed examples, resulting in unsatisfactory generalization on the validation set [Fig. 1(g)]. To address the instability of the random mix-up operation, our previous conference version, UMix (Shen et al., 2023) [Fig. 1(c)] synthesizes new samples with more high-confidence regions, enabling the model to produce high-quality predictions for these new samples. As shown in Fig. 1(f-g), the model trained with UMix samples demonstrates smoother and lower training losses but achieves inferior validation and test performance due to the limited effectiveness of consistency regularization. Furthermore, we investigated the inverse UMix (I-UMix) strategy [Fig. 1(d)]. Since this operation yields new samples with overly strong perturbations, the model shows severe oscillation in the training loss and obtains poor validation performance [Fig. 1(f-g)]. It can be observed that random perturbations tend to be uncontrollable, while mix-up operations with fixed rules result in trivial or overly strong perturbations, both of which limit the effectiveness of consistency regularization. Therefore, developing an adaptive perturbation/augmentation strategy for semi-supervised medical image segmentation is highly desirable.
@@
Conceptually, a model’s performance should vary dynamically with the number of training epochs increasing, thereby an appropriate mix-up operation should also dynamically adjust the perturbation degree based on the state of the segmentation model’s capability. Given this assumption, we propose Adaptive Mix (AdaMix), a novel dynamical and learnable perturbation algorithm that adaptively performs image mix-up based on a model’s learning state, for better achieving consistency regularization in semi-supervised medical image segmentation. Our key idea lies in introducing the concept of self-paced learning (Kumar et al., 2010; Jiang et al., 2014; Ma et al., 2017) into the design of the mix-up algorithm and providing more reasonable perturbed images for strong-weak pseudo supervision with gradually increasing sample complexity, as depicted in Fig. 1(e). The mix-up operation with varying degrees of perturbation leads to produce perturbed images with different levels of sample complexity. An image that contains more high-confidence regions is considered a simpler sample by a segmentation model, as it tends to produce more accurate predictions from such an image. Mixing with more high-confidence patches results in simpler perturbed images (with a weaker perturbation strength) and vice versa. More specifically, based on the training loss, AdaMix generates 1) a self-paced mask for determining whether an image is mixed with high-confidence patches or low-confidence ones, and simultaneously provides 2) a self-paced weight for determining the number of mix-up patches. One can observe from Fig. 1(f-g) that, compared with the random and fixed counterparts (i.e., CutMix (Yun et al., 2019) and UMix (Shen et al., 2023)), AdaMix is capable of maintaining stable converging ability during the training process and achieving better generalization in the validation and test phase. We empirically demonstrate that AdaMix can be seamlessly integrated into self-training, mean-teacher, and co-training (i.e., FixMatch (Sohn et al., 2020), MT (Tarvainen and Valpola, 2017), and CPS (Chen et al., 2021)) SSL frameworks, resulting in AdaMix-ST, AdaMix-MT, and AdaMix-CT frameworks, respectively, which achieve state-of-the-art performance in semi-supervised medical image segmentation. We evaluated the proposed frameworks on three public medical image segmentation datasets, involving 2D and 3D modalities. The results demonstrate the effectiveness of each component of our method and its superiority over the state-of-the-art. For instance, AdaMix-CT obtains 88.19% and 89.83% in terms of Dice Similarity Coefficient (DSC), outperforming MC-Net+ (Wu et al., 2022) by 2.09 % and 1.32% on the ACDC datasets under 10% and 20% labeled data, respectively.
@@
    New Problem: We investigate the problem of existing mix-up methods: they tend to generate perturbed images with uncontrollable, trivial, or overly strong perturbation degrees, which seriously affect consistency regularization and result in imprecise decision boundaries, particularly in heterogeneous lesion segmentation scenarios.
@@
    New method: We propose a novel algorithm, AdaMix, for image mix-up in a self-paced manner. To the best of our knowledge, this is the first work to adaptively perform mix perturbation based on a model’s state, where the segmentation model and the mix module collaboratively learn from each other. Furthermore, AdaMix is highly flexible and can be seamlessly integrated into self-training, mean-teacher, and co-training SSL paradigms.
@@
    New findings: We provide new insights: incorporating high-confidence patches into images generates simpler samples, which benefits the segmentation model during the initial training stage. In contrast, the low-confidence regions, which typically appear at segmentation boundaries, force the segmentation model to focus on these regions and enhance the model’s discriminative capability when mixed into images in the later training stages. We also found that, for consistency regularization, perturbation strategies are more important than learning paradigms (e.g., self-training, mean-teacher, and co-training).
@@ 2.1 Semi-Supervised Learning
Semi-supervised learning (SSL) is an effective learning strategy that addresses the issue of limited labeled data by utilizing a large amount of unlabeled data (Van Engelen and Hoos, 2020). Based on the design of the regularization term to leverage unlabeled data, many SSL algorithms have been proposed, such as adversarial learning (Springenberg, 2015; Miyato et al., 2018), pseudo labeling (Lee et al., 2013), consistency regularization (Rasmus et al., 2015; Laine and Aila, 2016). Recently, state-of-the-art SSL approaches have incorporated the idea of strong-weak pseudo supervision (Sohn et al., 2020) as the primary strategy of their frameworks. According to the learning paradigms, these approaches typically fall into three branches: self-training-based (Berthelot et al., 2019b, a; Sohn et al., 2020; Zhang et al., 2021; Wang et al., 2023a; Chen et al., 2023), mean-teacher-based (Tarvainen and Valpola, 2017) and co-training-based frameworks (Ke et al., 2019; Li et al., 2023). Similar to semi-supervised image classification tasks, the line of semi-supervised segmentation methods also can be divided into three branches: self-training-based (Zou et al., 2021; Yuan et al., 2021; Yang et al., 2022, 2023), mean-teacher-based (Yu et al., 2019; Hu et al., 2021; Xu et al., 2022; Liu et al., 2022), and co-training-based methods (Chen et al., 2021; Wang et al., 2023c; Luo et al., 2021a; Wu et al., 2021, 2022; Wang et al., 2023b; Shen et al., 2023). The self-training paradigm involves a single model that generates pseudo labels for training the model per se (Sohn et al., 2020). Mean-Teacher (Tarvainen and Valpola, 2017) can be considered an intermediate approach between self-training and co-training. It involves a teacher and a student model, where pseudo-supervision is conducted from the teacher to the student. A representative co-training paradigm, presented in CPS (Chen et al., 2021), comprises two (student) models that are cross-supervised by each other. We found that perturbation strategies on unlabeled data are more important than a specific learning paradigm. We propose a novel image mix-up algorithm that can be seamlessly applied to these paradigms and achieves state-of-the-art performance in semi-supervised medical image segmentation.
@@ 2.2 Mix-up Augmentation
Mix-up operations can be regarded as a line of data augmentation approaches in the form of blending two images and their corresponding labels to synthesize a new example. A series of mix-up strategies were initially proposed in supervised learning to generate new samples via randomly mixing images in image-level (Zhang et al., 2017), feature-level (Verma et al., 2019), and region-level (DeVries and Taylor, 2017; Yun et al., 2019). Later, additional information, such as saliency maps (Kim et al., 2020), was incorporated into the mix-up algorithms. Recently, image mix-up has been applied to perform perturbations for consistency regularization in a fully random manner (French et al., 2019; Chen et al., 2021; Verma et al., 2022) or according to predefined rules, e.g., replacing low-confidence patches with high-confidence ones guided by predictive uncertainty/confidence (Shen et al., 2023; Chi et al., 2024). However, the degree of random perturbations is uncontrollable, leading to overfitting to strongly perturbed samples, while the degree of fixed mix-up operations is trivial, both of which limit the effectiveness of consistency regularization. To solve this issue, we propose AdaMix which adaptively performs image mix-up perturbation in a self-paced learning manner for consistency regularization in semi-supervised medical image segmentation.

### v2

@@ Abstract
Mix-up is a key technique for consistency regularization-based semi-supervised learning methods, blending two or more images to generate strong-perturbed samples for strong-weak pseudo supervision. Existing mix-up operations are performed either randomly or with predefined fixed rules, such as replacing low-confidence patches with high-confidence ones. The former lacks control over the perturbation degree, leading to overfitting on randomly perturbed samples, while the latter tends to generate images with trivial perturbations, both of which limit the effectiveness of consistency regularization. This paper aims to answer the following question: How can image mix-up perturbation be adaptively performed during training? To this end, we propose an Adaptive Mix algorithm (AdaMix) for image mix-up in a self-paced learning manner. Given that, in general, a model’s performance gradually improves during training, AdaMix is equipped with a self-paced curriculum that, in the initial training stage, provides relatively simple perturbed samples and then gradually increases the difficulty of perturbed images by adaptively controlling the perturbation degree based on the model’s learning state estimated by a self-paced regularizer. We develop three frameworks with our AdaMix, i.e., AdaMix-ST, AdaMix-MT, and AdaMix-CT, for semi-supervised medical image segmentation. Extensive experiments on three public datasets, including both 2D and 3D modalities, show that the proposed frameworks are capable of achieving superior performance. For example, compared with the state-of-the-art, AdaMix-CT achieves relative improvements of 2.62% in Dice similarity coefficient and 48.25% in average surface distance on the ACDC dataset with 10% labeled data. The results demonstrate that mix-up operations with dynamically adjusted perturbation strength based on the segmentation model’s state can significantly enhance the effectiveness of consistency regularization. The code will be released at https://github.com/Senyh/AdaMix.
@@ 1 Introduction
Figure 1: Comparison among CutMix, UMix (Shen et al., 2023), inverse UMix, and our AdaMix. (a) Original and auxiliary images; (b) CutMix randomly replaces patches of the original image with patches from the auxiliary image, with an uncontrollable perturbation degree; (c) UMix replaces low-confidence patches in the original image with high-confidence ones from the auxiliary image, often generating perturbed examples with trivial perturbations; (d) I-UMix mixes high-confidence patches in the original image with low-confidence ones from the auxiliary image, yielding overly strong-perturbed images; (e) Our AdaMix synthesizes images with more high-confidence regions in the initial training stage and then gradually increases the perturbation degree to enhance the complexity of the perturbed images as training progresses; (f) The unsupervised learning loss curves; (g) The validation loss curves (and the Dice Similarity Coefficients on the ACDC test set).
Medical image segmentation for delineating tissues, organs, and lesions from different imaging modalities is critical for computer-aided clinical applications, such as prognosis and preoperative planning. Deep learning has greatly advanced the progress of medical image segmentation, relying on its powerful representation capability and large amounts of labeled data (Ronneberger et al., 2015; Milletari et al., 2016; Chen et al., 2021a; Wang et al., 2022; Wu et al., 2024a). However, obtaining pixel-level annotations is laborious and expensive in medical images, as it requires expert knowledge. Semi-supervised learning (SSL) has the potential to alleviate the label scarcity problem by leveraging a large amount of unlabeled data (Van Engelen and Hoos, 2020).
@@
Consistency regularization, one of the most widely used SSL techniques, encourages model predictions to be invariant to input perturbations, aiming to explore supervision information from unlabeled data (Rasmus et al., 2015; Laine and Aila, 2016; Tarvainen and Valpola, 2017). Building upon the idea of consistency regularization, strong-weak pseudo supervision (Sohn et al., 2020) enforces consistency between the predicted pseudo labels of strongly and weakly perturbed unlabeled images, thereby sufficiently extracting supervisory signals from unlabeled data (Sohn et al., 2020; Chen et al., 2023, 2021b; Wu et al., 2022; Shen et al., 2023; Wang et al., 2023b). The key challenge of strong-weak pseudo supervision is determining appropriate strong perturbation strategies for generating strong-perturbed images. Improper perturbations cannot match the training model’s ability to capture supervision signals from unlabeled data, misleading the training process (Shen et al., 2023). Recently, CutMix [^1] CutMix is formulated as $\tilde{x}=x_{o}\times m+x_{a}\times(1-m)$ and $\tilde{y}=y_{o}\times m+y_{a}\times(1-m)$, where $x_{o}$ / $x_{a}$ denotes an original/auxiliary image, $y_{o}$ / $y_{a}$ refers to the corresponding (pseudo) label, and $m$ is a randomly generated mask. $\tilde{x}$ / $\tilde{y}$ represents the mixed image / (pseudo) label.(Yun et al., 2019), incorporated with image color space transformation, has been widely used to form the strong perturbations (French et al., 2019; Chen et al., 2021b; Yang et al., 2022, 2023a). As illustrated in Fig. 1(b), CutMix randomly mixes image patches between an original image and an auxiliary image to synthesize a strong-perturbed example. However, such a random mix operation with an uncontrollable perturbation degree leads to an oscillating and unstable consistency regularization process [Fig. 1(f)], where the model is prone to over-fit to the randomly perturbed examples, resulting in unsatisfactory generalization on the validation set [Fig. 1(g)]. To address the instability of the random mix-up operation, our previous work, UMix (Shen et al., 2023) [Fig. 1(c)] synthesizes strong-perturbed images with more high-confidence regions, enabling the model to produce high-quality predictions for these samples. As shown in Fig. 1(f-g), the model trained with UMix images demonstrates smoother and lower training losses but achieves inferior validation and test performance due to the limited effectiveness of consistency regularization. Furthermore, we investigated the inverse UMix (I-UMix) strategy [Fig. 1(d)]. Since this operation yields new samples with overly strong perturbations, the model shows severe oscillation in the training loss and obtains poor validation performance [Fig. 1(f-g)]. It can be observed that random perturbations tend to be uncontrollable, while mix-up operations with fixed rules result in trivial or over-strong perturbations, both of which limit the effectiveness of consistency regularization. Therefore, developing an adaptive perturbation/augmentation strategy for semi-supervised medical image segmentation is highly desirable.
@@
Conceptually, a model’s performance should vary dynamically as the number of training epochs increases; thereby, an appropriate mix-up operation should also dynamically adjust the perturbation degree based on the state of the segmentation model’s capability. Given this assumption, we propose Adaptive Mix (AdaMix), a novel dynamical and learnable perturbation algorithm that adaptively performs image mix-up based on a model’s learning state, for better achieving consistency regularization in semi-supervised medical image segmentation. Our key idea lies in introducing the concept of self-paced learning (Kumar et al., 2010; Jiang et al., 2014; Ma et al., 2017) into the design of the mix-up algorithm and providing more reasonable perturbed images with gradually increasing sample complexity for strong-weak pseudo supervision, as depicted in Fig. 1(e). The mix-up operation with varying degrees of perturbation leads to the generation of perturbed images with different levels of sample complexity. An image that contains more high-confidence regions is considered an easier sample by a segmentation model, as it tends to produce more accurate predictions from such an image; in contrast, mixing with more high-confidence patches results in a harder example (with a stronger perturbation strength). More specifically, based on the training loss, AdaMix generates 1) a self-paced mask for determining whether an image is mixed with high-confidence patches or low-confidence ones, and simultaneously provides 2) a self-paced weight for determining the number of mix-up patches. One can observe from Fig. 1(f-g) that, compared with the random and fixed counterparts (i.e., CutMix (Yun et al., 2019) and UMix (Shen et al., 2023)), AdaMix is capable of maintaining stable converging ability during the training process and achieving better generalization in the validation and test phase. We empirically demonstrate that AdaMix can be seamlessly integrated into self-training, mean-teacher, and co-training (i.e., FixMatch (Sohn et al., 2020), MT (Tarvainen and Valpola, 2017), and CPS (Chen et al., 2021b)) SSL paradigms, resulting in AdaMix-ST, AdaMix-MT, and AdaMix-CT frameworks, respectively, which achieve state-of-the-art performance in semi-supervised medical image segmentation tasks. We evaluated the proposed frameworks on three public medical image segmentation datasets, involving 2D and 3D modalities. The results demonstrate the effectiveness of each component of our method and its superiority over the state-of-the-art.
@@
    New Perspective: We identify the limitations in existing mix-up methods from a dynamic model training perspective: they perform image mix-up either in a random manner or based on pre-defined fixed rules, resulting in perturbed images with uncontrollable, trivial, or over-strong perturbation degrees, which seriously undermine the effectiveness of consistency regularization. We suggest that an adaptive mix-up scheme with awareness of the model’s learning state is critical for consistency regularization.
@@
    New method: We propose a novel algorithm, AdaMix, for image mix-up in a self-paced manner. To the best of our knowledge, this is the first work to adaptively perform mix-up perturbation based on a model’s state, where the segmentation model and the AdaMix algorithm collaboratively learn from each other. Furthermore, AdaMix is highly flexible and can be seamlessly integrated into self-training, mean-teacher, and co-training SSL paradigms.
@@
    New findings: We provide new insights: incorporating high-confidence patches into images generates simpler samples, which benefits the segmentation model during the initial training stage. In contrast, the low-confidence regions, which typically appear at segmentation boundaries, force the segmentation model to focus on these hard regions and enhance the model’s discriminative capability when mixed into images in the later training stages. We also reveal that perturbation strategies are more important than learning paradigms for consistency regularization.
@@ 2.1 Semi-Supervised Learning
Semi-supervised learning (SSL) is an effective learning strategy that addresses the issue of limited labeled data by utilizing a large amount of unlabeled images (Van Engelen and Hoos, 2020). Based on the design of the regularization term to leverage unlabeled data, many SSL algorithms have been proposed, such as adversarial learning (Springenberg, 2015; Miyato et al., 2018), pseudo labeling (Lee et al., 2013), and consistency regularization (Rasmus et al., 2015; Laine and Aila, 2016). Recently, state-of-the-art SSL approaches have incorporated the idea of strong-weak pseudo supervision (Sohn et al., 2020) as the primary strategy of their frameworks. According to the learning paradigms, these approaches typically fall into three branches: self-training-based (Berthelot et al., 2019b, a; Sohn et al., 2020; Zhang et al., 2021; Wang et al., 2023a; Chen et al., 2023), mean-teacher-based (Tarvainen and Valpola, 2017) and co-training-based frameworks (Ke et al., 2019; Li et al., 2023).
Similar to semi-supervised image classification tasks, the line of semi-supervised segmentation methods also can be divided into three branches: self-training-based (Zou et al., 2021; Yuan et al., 2021; Yang et al., 2022, 2023a), mean-teacher-based (Yu et al., 2019; Hu et al., 2021; Xu et al., 2022; Liu et al., 2022a), and co-training-based methods (Chen et al., 2021b; Wang et al., 2023c; Luo et al., 2021a; Wu et al., 2021, 2022; Wang et al., 2023b; Shen et al., 2023). The self-training paradigm involves a single model that generates pseudo labels for training the model per se (Sohn et al., 2020). Mean-Teacher (Tarvainen and Valpola, 2017) can be considered an intermediate approach between self-training and co-training. It involves a teacher and a student model, where pseudo-supervision is conducted from the teacher to the student. A representative co-training paradigm, presented in CPS (Chen et al., 2021b), comprises two (student) models that are cross-supervised by each other. We find that perturbation strategies are more important than a specific learning paradigm. We propose a novel image mix-up algorithm that can be seamlessly applied to these paradigms and achieves state-of-the-art performance in semi-supervised medical image segmentation.
@@ 2.2 Mix-up Augmentation
Mix-up operations can be regarded as a line of data augmentation approaches that blend two or more images and their corresponding labels to synthesize a new example. A series of mix-up strategies were initially proposed in supervised learning to generate new samples via randomly mixing images in image-level (Zhang et al., 2017), feature-level (Verma et al., 2019), and region-level (DeVries and Taylor, 2017; Yun et al., 2019; Liu et al., 2022b). Furthermore, additional information, such as saliency maps (Kim et al., 2020, 2021; Uddin et al., 2021; Kang and Kim, 2023), was incorporated into the mix-up algorithms. These methods extend training data distributions and impose strong regularization during model training, implicitly controlling model complexity and enhancing model robustness.
Recently, mix-up algorithms have been utilized to introduce strong perturbations for consistency regularization in semi-supervised learning frameworks for image recognition (Berthelot et al., 2019b, a; Sohn et al., 2020; Verma et al., 2022) and semantic segmentation (French et al., 2019; Zou et al., 2020; Chen et al., 2021b; Yang et al., 2023b; Shen et al., 2023; Chi et al., 2024). Specifically, they are performed in fully random manners (e.g., Mixup, Cutout, or CutMix)  (Berthelot et al., 2019b, a; Sohn et al., 2020; French et al., 2019; Zou et al., 2020; Chen et al., 2021b; Verma et al., 2022), or according to predefined fixed rules (e.g., UMix that replaces low-confidence patches with high-confidence ones guided by predictive uncertainty/confidence) (Shen et al., 2023; Chi et al., 2024). However, the strength of the random perturbation is uncontrollable, leading to overfitting to randomly perturbed samples, while the degree of the fixed mix-up operations is trivial or over-strong. Both cases generate improper perturbed images limiting the effectiveness of consistency regularization. To address this issue, an adaptive mix-up perturbation scheme is highly desirable for effective consistency regularization in semi-supervised medical image segmentation.
Figure 2: Overview of the proposed Adaptive Mix framework for semi-supervised medical image segmentation. It includes our Adaptive Mix (AdaMix) algorithm for generating perturbed images and a semi-supervised learning paradigm for providing pseudo labels and conducting consistency regularization. AdaMix can be seamlessly applied to the self-training, mean-teacher, and co-training paradigms, resulting in AdaMix-ST, AdaMix-MT, and AdaMix-CT frameworks, respectively.
2.3 Medical Image Segmentation
Deep learning-based methods have achieved remarkable progress in medical image segmentation tasks (Ronneberger et al., 2015; Çiçek et al., 2016; Milletari et al., 2016). Specifically, building upon fully convolutional networks (Long et al., 2015), U-Net and its variants (Ronneberger et al., 2015; Çiçek et al., 2016; Milletari et al., 2016; Zhou et al., 2018; Oktay et al., 2018; Isensee et al., 2021) have been the mainstream baseline segmentation models widely adopted in further studies, such as semi-supervised segmentation and related tasks (Yu et al., 2019; Wu et al., 2021; Shen et al., 2023). Furthermore, to address the issues of multi-scale heterogeneous organs and lesions in medical imaging, numerous transformer-based models  (Chen et al., 2021a; Valanarasu et al., 2021; Cao et al., 2022; Wang et al., 2022, 2024) have been applied to model long-range dependencies among segmentation targets. Meanwhile, diffusion models-based approaches have recently been developed to improve the robustness of segmentation models against noise and artifacts of medical images (Wu et al., 2024a, b). More recently, Segment Anything Models (Ma et al., 2024; Wei et al., 2024) emerged as a foundation model aiming at universal medical image segmentation across a wide spectrum of tasks. Although these existing methods show promising results, their performance relies heavily on large-scale datasets with numerous pixel-level annotations, which are time-consuming and expertise-demanding in specific clinical applications, e.g. assessment of cardiac ejection fraction and diagnosis of skin melanoma. In contrast, our AdaMix algorithm can function as a plug-and-play, model-agnostic module, enabling consistency regularization on these fully supervised models to leverage unlabeled data for alleviating the annotation burden.